# ATK: Automatic Task-driven Keypoint Selection for Robust Policy Learning

### v1

@@ 2 Problem Formulation
Imitation Learning. The imitation learning setting is provided with an offline dataset of expert behavior data in lieu of a simulator. The dataset $\mathcal{D}=\{(o_{i},a_{i})\}_{i=1}^{N}$ is drawn from an expert – $\pi^{*}$. Imitation learning settings both train and test in the real world from raw camera observations. However, for imitation learning to be visually “robust," we consider training on a setting with one emission function, $f_{\text{train}}(s)$, and evaluating with another emission function, $f_{\text{test}}(s)$. We can formulate this via a simple dual MDP formulations: $\mathcal{M}_{\text{train}}=(\mathcal{O}_{\text{train}},\mathcal{S},\mathcal{A},\mathcal{P},\rho_{0},\mathcal{R},\gamma)$ for training and $\mathcal{M}_{\text{test}}=(\mathcal{O}_{\text{real}},\mathcal{A},\mathcal{P},\rho_{0},\mathcal{R},\gamma)$ for evaluation. The same underlying state $s$ goes through different emission functions and leads to different observations at training, $o_{\text{train}}=f_{\text{train}}(s)$, and testing, $o_{\text{test}}=f_{\text{test}}(s)$. For instance, this approach could involve training for robotic manipulation on a clear tabletop and then testing with changes in background, lighting or distractor objects. The challenge lies in the mismatch between $\mathcal{O}_{\text{train}}$ and $\mathcal{O}_{\text{test}}$.
@@ 3.1 Keypoints as Policy Representations
Keypoints, a widely used representation in computer vision, are salient locations in an image that are useful for identifying and describing objects or features. Formally, a keypoint is defined as a specific position $k_{t}^{i}=(x_{t}^{i},y_{t}^{i})$ in the 2D image plane at time $t$. A set of $N$ keypoints, $\{k_{t}^{i}\}_{i=1}^{N}$, provides a compact scene representation. The number and selection of keypoints can be dynamically adjusted based on task complexity and requirements. However, what makes keypoints impactful for robust policy learning is the proliferation of robust tracking algorithms  [16, 15], trained on web-scale data, that maintain dense correspondences across frames despite visual scene-level and instance-level variations. We bring this robustness to bear on policy learning. To track keypoints over time, we initialize keypoints $\{k_{t}^{i}\}_{i=1}^{N}$ at $t=0$ and then use tracking methods [21, 16, 15] to maintain robust semantic correspondences of these points across time steps. We formalize tracking as a correspondence function $h_{\mathcal{C}}$ that updates keypoint locations at each time step while providing correspondence measurement scores.
Given a set of initial keypoint positions, $\{k_{0}^{i}\}_{i=1}^{N}=(x_{0}^{i},y_{0}^{i},...,x_{n}^{i},y_{n}^{i})$, a particular set of their positions at time $t$ is updated as: $\{k_{t}^{i}\}_{i=1}^{N}=h_{\mathcal{C}}(\{k_{t-1}^{i}\}_{i=1}^{N},o_{t})$. The observation at time t, $o_{t}$, is an (ordered) set of $N$ keypoints $\{k_{t}^{i}\}_{i=1}^{N}$ that is updated iteratively through the correspondence function $h_{\mathcal{C}}$. The keypoint locations $k_{t}^{i}=(x_{t}^{i},y_{t}^{i})$ correspond to the current planar positions of points that semantically correspond to the initially chosen points $\{k_{0}^{i}\}_{i=1}^{N}$.
The core challenges in leveraging keypoints as a policy representation are the selection of initial keypoints $\{k_{0}^{i}\}_{i=1}^{N}$, the tracking of them through time, and the transferring of resulting keypoint-based policies across widely varying deployment scenarios. The choice of keypoints is inherently task-specific since different tasks require focusing on distinct elements in the scene. For example, in the kitchen scene shown in Fig 1, the keypoints on the blanket are crucial for the blanket-hanging task, whereas the pan-placement and grasping tasks require keypoints on both the pan and other objects.
@@ 3.2 Automatic Task-Driven Keypoint Selection: Training
Our work selects a minimal set of task-relevant keypoints as a representation to enable robust policy transfer. Formally, we aim to identify a minimal set of $N$ task-relevant keypoints, $\{k_{0}^{i}\}_{i=1}^{N}$, that enable training a near-optimal policy while being easily tracked with $h_{\mathcal{C}}$. Our keypoint selection is based on two criteria: (1) realizability of the optimal policy, i.e., the selected keypoints must capture all necessary information to learn a near optimal policy for the task, and (2) trackability, i.e., the chosen keypoints must be reliably and consistently trackable using an available correspondence function $h_{\mathcal{C}}$. Realizability and trackability are naturally interconnected concepts; representations that are not trackable are not consistent, making it impossible to realize an optimal policy. Given a larger set of $M$ candidate keypoints $\{k_{t}^{i}\}_{i=1}^{M}$, how do we select a minimal set of $N$ keypoints, with $N\ll M$, to satisfy realizability and trackability?
@@
To ensure that each keypoint remains semantically and spatially consistent across trajectories, we first identify a single image frame that captures the complete task context for solving the task, denoted as the canonical template $I_{\text{can}}$. We then randomly sample $C$ candidate keypoints $\{k_{j}^{c}\}_{j=1}^{C}$ on the canonical image $I_{\mathrm{can}}$. Given an expert dataset $\mathcal{D}=\bigl{\{}(o_{i},a_{i})\bigr{\}}_{i=1}^{M},$ we use the correspondence function $h_{\mathcal{C}}$ to align and propagate these candidates across all trajectories and timesteps, producing a keypoint-annotated dataset $\mathcal{D}_{k}=\bigl{\{}(o_{i,t},\,K_{i,t},\,a_{i})\bigr{\}}_{i=1,t=1}^{M,T_{i}},$ where $K_{i,t}$ is the set of all tracked keypoint positions for observation $o_{i,t}$. From $\mathcal{D}_{k}$, we jointly learn a sparse masking model $\mathbb{M}_{\phi}$ and a downstream keypoint-based policy $\pi^{k}_{\theta}$. As shown in Fig. 2, the $M$ candidate keypoints are fed into $\mathbb{M}_{\phi}$, which outputs $M$ independent Bernoulli probabilities, with each representing the likelihood of retaining the corresponding keypoint from the input candidate set. Sampling a binary mask $m\in\{0,1\}^{M}$ then zeroes out the unselected keypoints, yielding a reduced set $\widetilde{K}$. Finally, $\widetilde{K}$ is passed to $\pi^{k}_{\theta}$, which produces the action distribution, as is standard in imitation learning. This unified architecture first applies pointwise masking to select the minimal keypoint subset and then predicts actions from those selected points.
To select the minimal set of $N$ keypoints, we enforce a sparse information bottleneck on the mask:
$\min-\mathbb{E}_{(k,a)\sim\mathcal{D}}\left[\log\pi^{k}_{\theta}(a^{*}_{t}\mid\{k_{t}^{i}\}_{i=1}^{N})\right]+\alpha\|\mathbb{M}_{\phi}(\{k_{t}^{i}\}_{i=1}^{N})\|_{1}.$
Intuitively, this training procedure filters out points that are (1) irrelevant to predicting optimal actions and (2) challenging to track using $h_{\mathcal{C}}$ since their representations over time ($(\{k_{t}^{i}\}_{i=1}^{M})$ are unreliable markers of optimal actions. Since the mask sampling is discrete and non-differentiable, we employ the Gumbel-softmax relaxation [22] to enable gradient-based optimization.
@@ 3.3 Inference with Task-Driven Keypoints
The preceding procedure learns a masking model $q_{\phi}(m\mid K)$ and a policy $\pi_{\theta}(a\mid m\odot K)$. How can we use them for robust policy inference at test time? Whether it be a transfer from simulation to reality or from one imitation learning scenario at training time to another at test time, the inference procedure remains the same. For inference, it is important to transfer the minimal set of keypoints selected at training time to a variety of visually diverse test time scenarios. Since tracking of keypoints is performed by robust web-scale visual trackers [16, 15], once the initial selected set of keypoints is identified in the real world at test time, subsequent tracking is not affected by the visual gap. This lets us focus solely on transferring the initial set of keypoints.
@@
Assuming the correspondence function $h_{\mathcal{C}}$ provides confidence scores for each candidate pair, at test time we select the initial-image points with the highest scores relative to the training canonical keypoints $\{k_{0}^{i}\}_{i=1}^{N}$, ensuring accurate matches. These transferred keypoints are then tracked via $h_{\mathcal{C}}$ and used directly as input to the deployed policy $\pi_{\theta}$ without requiring additional masking at test time. Though many implementations of $h_{\mathcal{C}}$ are possible, we employ diffusion-model features [17, 23] for their strong visual robustness and matching reliability. We provide the pseudocode for applying $ATK$ on new task in Algorithm 1 and recommended hyperparameter selection in Table 6.
Mask model $\mathbb{M}_{\phi}$, keypoint policy $\pi_{\theta}$
// Initialize keypoints on canonical frame
Choose canonical image $I_{\text{can}}$ capturing full task context
Sample $C$ candidate keypoints $K_{\text{can}}\!=\!\{k_{j}^{c}\}_{j=1}^{C}$ on $I_{\text{can}}$
// Propagate & label entire dataset
for $i\leftarrow 0$ to $M-1$ do $\triangleright$ for all trajectories
     for $t\leftarrow 0$ to $T_{i}-1$ do
         if $t=0$ then
              $K_{i,0}\leftarrow h_{\mathcal{C}}\!\bigl{(}K_{\text{can}},\,o_{i,0},\,I_{\text{can}}\bigr{)}$triangleright initial alignment to canonical frame
         else
              $K_{i,t}\leftarrow h_{\mathcal{C}}\!\bigl{(}K_{i,t-1},\,o_{i,t}\bigr{)}$triangleright track from previous frame
         end if
     end for
end for
Construct $\mathcal{D}_{k}=\{(o_{i,t},K_{i,t},a_{i})\}$
//Training
while not converged do
     Sample mini-batch $\mathcal{B}\subset\mathcal{D}_{k}$
     Draw mask $m\sim\mathbb{M}_{\phi}(K)$triangleright Gumbel–Softmax
     $\widetilde{K}\leftarrow m\odot K$triangleright retain selected points
     Update $\phi,\theta$ with Adam to minimise $\mathcal{L}$
end while
return $(\mathbb{M}_{\phi},\pi_{\theta})$
function Evaluation
     // Transfer keypoints to test scene
     $K_{0}\leftarrow h_{\mathcal{C}}\!\bigl{(}K_{\text{can}},\,I_{\text{can}},\,O_{init}\bigr{)}$triangleright highest-score matches
     Track $K_{t}=h_{\mathcal{C}}(K_{t-1},o_{t})$ for $t\geq 1$
     return $\pi_{\theta}(K_{t})$ at each step
end function
@@ 4.2 Sim-to-Real Transfer of Keypoint-based Policies
$ATK$ transfers from sim-to-real demonstrating visual robustness. As shown in Fig 5, keypoint-based policies maintain high success rates in the real world compared to alternative modalities, showcasing strong resilience against randomized object poses or background variations. We provide aggregate performance across different distractors, providing a detailed per-disturbance breakdown in the Appendix. E. Although extreme distractions, such as flashing light or occlusions, can disrupt tracking and decrease performance, ATKconsistently outperforms RGB, depth, and point-cloud based policies at transfer. The gap is worth noting in tasks involving transparent objects (e.g., glass) and fine-grained manipulation (e.g., clock tasks).
@@ 4.3 Robust Imitation Learning with Keypoint-based Policies
4.4 Qualitative Visualization of Keypoints Learned by $ATK$
$ATK$ chooses interpretable and task-relevant keypoints. In Fig 8, we show the keypoints $ATK$ selects, focusing on task-relevant parts. Corresponding visualizations for imitation experiments are done in Fig 1.The chosen keypoints correspond to semantically meaningful elements of the scene. In multifunctional cases (e.g., kitchen, clock), they are specialized to the utility. We see that chosen keypoints transfer accurately from simulation to the real world and from training to testing. The selected keypoints are resilient to visual variations in the scene, including distractors, lighting changes, and background changes.
@@ 5 Related Work
Keypoints as Representations for Learning-based Control. Keypoints have been utilized as robust state representations for robotic manipulation in several prior works [36, 37, 38, 39]. They have been applied in areas including deformable object manipulation [40, 41], few-shot imitation learning [37], model-based reinforcement learning [39], and learning from videos [36, 42]. However, these approaches often rely on heuristic or manual keypoint selection [43, 44]. Our work differs by introducing a task-driven method for automatic keypoint selection. The procedure produces flexible but expressive representations that generalize across rigid, deformable, and transparent objects, and—crucially—are robust to challenging visual disturbances.
@@ 6 Limitations and Conclusion
We present ATK, a system for automatically selecting task-relevant keypoints, learning keypoint-based policies from these representations, and successfully transferring them to the real-world evaluation scenario. Though promising, the system faces challenges in tracking and optimization. The use of 2D keypoints makes the policy sensitive to camera perspective changes, and off-the-shelf tracking modules may lack robustness for robotic applications with uncommon visual data. Additionally, the method is sensitive to hyperparameters due to the non-smooth nature of the optimization problem, making tuning difficult. Nevertheless, our work demonstrates the robustness of keypoint-based policies and provides an effective approach for automatic keypoint selection. Developing more automated and robust techniques to address these challenges would further extend its applicability.
@@ Acknowledgments
This work was (partially) funded by grants from the National Science Foundation NRI (#2132848), DARPA RACER (#HR0011-21-C-0171), the Office of Naval Research (#N00014-24-S-B001 and #2022-016-01 UW), and funding from the Toyota Research Institute through the University Research Program 3.0. We gratefully acknowledge gifts from Amazon, Collaborative Robotics, Cruise, and others. We would also like to thank Emma Romig, Grant Shogren for help in setting up the hardware environments for this effort. We thank members of the Washington Embodied Intelligence and Robotics Development Lab and the Personal Robotics Lab for their thoughtful comments and feedback on versions of this draft.

### v2

@@ 2 Problem Formulation
Imitation Learning. The imitation learning setting is provided with an offline dataset of expert behavior data in lieu of a simulator. The dataset $\mathcal{D}=\{(o_{i},a_{i})\}_{i=1}^{M}$ is drawn from an expert – $\pi^{*}$. Imitation learning settings both train and test in the real world from raw camera observations. However, for imitation learning to be visually “robust," we consider training on a setting with one emission function, $f_{\text{train}}(s)$, and evaluating with another emission function, $f_{\text{test}}(s)$. We can formulate this via a simple dual MDP formulations: $\mathcal{M}_{\text{train}}=(\mathcal{O}_{\text{train}},\mathcal{S},\mathcal{A},\mathcal{P},\rho_{0},\mathcal{R},\gamma)$ for training and $\mathcal{M}_{\text{test}}=(\mathcal{O}_{\text{real}},\mathcal{A},\mathcal{P},\rho_{0},\mathcal{R},\gamma)$ for evaluation. The same underlying state $s$ goes through different emission functions and leads to different observations at training, $o_{\text{train}}=f_{\text{train}}(s)$, and testing, $o_{\text{test}}=f_{\text{test}}(s)$. For instance, this approach could involve training for robotic manipulation on a clear tabletop and then testing with changes in background, lighting or distractor objects. The challenge lies in the mismatch between $\mathcal{O}_{\text{train}}$ and $\mathcal{O}_{\text{test}}$.
@@ 3.1 Keypoints as Policy Representations
Keypoints, a widely used representation in computer vision, are salient locations in an image that are useful for identifying and describing objects or features. Formally, a keypoint is defined as a specific position $k_{i}^{t}=(x_{i}^{t},y_{i}^{t})$ in the 2D image plane at time $t$. A set of $N$ keypoints, $\{k_{i}^{t}\}_{i=1}^{N}=(x_{1}^{t},y_{1}^{t},...,x_{N}^{t},y_{N}^{t})$, provides a compact scene representation. The number and selection of keypoints can be dynamically adjusted based on task complexity and requirements. However, what makes keypoints impactful for robust policy learning is the proliferation of robust tracking algorithms  [16, 15], trained on web-scale data, that maintain dense correspondences across frames despite visual scene-level and instance-level variations. We bring this robustness to bear on policy learning. To track keypoints over time, we initialize keypoints $\{k_{i}^{t}\}_{i=1}^{N}$ at $t=0$ and then use tracking methods [21, 16, 15] to maintain robust semantic correspondences of these points across time steps. We formalize tracking as a correspondence function $h_{\mathcal{C}}$ that updates keypoint locations at each time step while providing correspondence measurement scores.
Given a set of initial keypoint positions, $\{k_{i}^{t}\}_{i=1}^{N}=(x_{1}^{t},y_{1}^{t},...,x_{N}^{t},y_{N}^{t})$, a particular set of their positions at time $t$ is updated as: $\{k_{i}^{t}\}_{i=1}^{N}=h_{\mathcal{C}}(\{k_{i}^{t-1}\}_{i=1}^{N},I_{t})$, where $I_{t}$ is the current image. The observation at time t, $o_{t}$, is an (ordered) set of $N$ keypoints $\{k_{i}^{t}\}_{i=1}^{N}$ that is updated iteratively through the correspondence function $h_{\mathcal{C}}$. The keypoint locations $k_{i}^{t}=(x_{i}^{t},y_{i}^{t})$ correspond to the current planar positions of points that semantically correspond to the initially chosen points $\{k_{i}^{0}\}_{i=1}^{N}$.
The core challenges in leveraging keypoints as a policy representation are the selection of initial keypoints $\{k_{i}^{0}\}_{i=1}^{N}$, the tracking of them through time, and the transferring of resulting keypoint-based policies across widely varying deployment scenarios. The choice of keypoints is inherently task-specific since different tasks require focusing on distinct elements in the scene. For example, in the kitchen scene shown in Fig 1, the keypoints on the blanket are crucial for the blanket-hanging task, whereas the pan-placement and grasping tasks require keypoints on both the pan and other objects.
@@ 3.2 Automatic Task-Driven Keypoint Selection: Training
Our work selects a minimal set of task-relevant keypoints as a representation to enable robust policy transfer. Formally, we aim to identify a minimal set of $N$ task-relevant keypoints, $\{k_{i}\}_{i=1}^{N}$, that enable training a near-optimal policy while being easily tracked with $h_{\mathcal{C}}$. Our keypoint selection is based on two criteria: (1) realizability of the optimal policy, i.e., the selected keypoints must capture all necessary information to learn a near optimal policy for the task, and (2) trackability, i.e., the chosen keypoints must be reliably and consistently trackable using an available correspondence function $h_{\mathcal{C}}$. Realizability and trackability are naturally interconnected concepts; representations that are not trackable are not consistent, making it impossible to realize an optimal policy. Given the full set of $N$ candidate keypoints $\{k_{i}^{t}\}_{i=1}^{N}$, how do we select a minimal set of $K$ keypoints, with $K\ll N$, to satisfy realizability and trackability?
@@
To ensure that each keypoint remains semantically and spatially consistent across trajectories, we first identify a single image frame that captures the complete task context for solving the task, denoted as the canonical template $I_{\text{can}}$. We then randomly sample $C$ candidate keypoints $\{k_{j}^{0}\}_{j=1}^{C}$ on the canonical image $I_{\mathrm{can}}$. Given an expert dataset $\mathcal{D}=\bigl\{(o_{i},a_{i})\bigr\}_{i=1}^{M},$ we use the correspondence function $h_{\mathcal{C}}$ to align and propagate these candidates across all trajectories and timesteps, producing a keypoint-annotated dataset $\mathcal{D}_{k}=\bigl\{(o_{i,t},\,K_{i,t},\,a_{i})\bigr\}_{i=1,t=1}^{M,T_{i}},$ where $K_{i,t}$ is the set of all tracked keypoint positions for observation $o_{i,t}$. From $\mathcal{D}_{k}$, we jointly learn a sparse masking model $\mathbb{M}_{\phi}$ and a downstream keypoint-based policy $\pi^{k}_{\theta}$. As shown in Fig. 2, the $M$ candidate keypoints are fed into $\mathbb{M}_{\phi}$, which outputs $M$ independent Bernoulli probabilities, with each representing the likelihood of retaining the corresponding keypoint from the input candidate set. Sampling a binary mask $m\in\{0,1\}^{M}$ then zeroes out the unselected keypoints, yielding a reduced set $\widetilde{K}$. Finally, $\widetilde{K}$ is passed to $\pi^{k}_{\theta}$, which produces the action distribution, as is standard in imitation learning. This unified architecture first applies pointwise masking to select the minimal keypoint subset and then predicts actions from those selected points.
To select the minimal set of $K$ keypoints, we enforce a sparse information bottleneck on the mask:
$\min-\mathbb{E}_{(k,a)\sim\mathcal{D}}\log\pi_{\theta}^{k}\!\left(a_{t}^{*}\,\middle|\,\mathbb{M}_{\phi}\!\bigl(\{k_{i}^{t}\}_{i=1}^{N}\bigr)\right)+\alpha\|\mathbb{M}_{\phi}(\{k_{i}^{t}\}_{i=1}^{N})\|_{1}.$
Intuitively, this training procedure filters out points that are (1) irrelevant to predicting optimal actions and (2) challenging to track using $h_{\mathcal{C}}$ since their representations over time $(\{k_{i}^{t}\}_{i=1}^{N})$ are unreliable markers of optimal actions. Since the mask sampling is discrete and non-differentiable, we employ the Gumbel-softmax relaxation [22] to enable gradient-based optimization.
@@ 3.3 Inference with Task-Driven Keypoints
The preceding procedure learns a masking model $\mathbb{M}_{\phi}$ and a policy $\pi_{\theta}^{k}\!\left(a_{t}^{*}\,\middle|\,\mathbb{M}_{\phi}\!\bigl(\{k_{i}^{t}\}_{i=1}^{N}\bigr)\right)$. How can we use them for robust policy inference at test time? Whether it be a transfer from simulation to reality or from one imitation learning scenario at training time to another at test time, the inference procedure remains the same. For inference, it is important to transfer the minimal set of keypoints selected at training time to a variety of visually diverse test time scenarios. Since tracking of keypoints is performed by robust web-scale visual trackers [16, 15], once the initial selected set of keypoints is identified in the real world at test time, subsequent tracking is not affected by the visual gap. This lets us focus solely on transferring the initial set of keypoints.
@@
Assuming the correspondence function $h_{\mathcal{C}}$ provides confidence scores for each candidate pair, at test time we select the initial‐image points with the highest scores relative to the training canonical keypoints $\{k_{i}\}_{i=1}^{N}$, ensuring accurate matches. These transferred keypoints are then tracked via $h_{\mathcal{C}}$ and used directly as input to the deployed policy $\pi_{\theta}$ without requiring additional masking at test time. Though many implementations of $h_{\mathcal{C}}$, we adopt the method using diffusion-feature  [17, 23] for initial set transfer from canonical image due to its strong visual robustness and reliable matching, and employ Cotracker [16] for subsequent tracking. We provide the pseudocode for applying $ATK$ on a new task in Algorithm 1 and recommended hyperparameter selection in Table 7.
@@ 4.2 Sim-to-Real Transfer of Keypoint-based Policies
ATKATK transfers from sim-to-real demonstrating visual robustness. As shown in Fig 5, keypoint-based policies maintain high success rates in the real world compared to alternative modalities, showcasing strong resilience against randomized object poses or background variations. We provide aggregate performance across different distractors, providing a detailed per-disturbance breakdown in the Appendix. E. Although extreme distractions, such as flashing light or occlusions, can disrupt tracking and decrease performance, ATKconsistently outperforms RGB, depth, and point-cloud based policies at transfer. The gap is worth noting in tasks involving transparent objects (e.g., glass) and fine-grained manipulation (e.g., clock tasks).
@@ 4.3 Robust Imitation Learning with Keypoint-based Policies
4.4 Qualitative Visualization of Keypoints Learned by ATKATK
ATKATK chooses interpretable and task-relevant keypoints. In Fig 7, we show the keypoints $ATK$ selects, focusing on task-relevant parts. Corresponding visualizations for imitation experiments are done in Fig 1.The chosen keypoints correspond to semantically meaningful elements of the scene. In multifunctional cases (e.g., kitchen, clock), they are specialized to the utility. We see that chosen keypoints transfer accurately from simulation to the real world and from training to testing. The selected keypoints are resilient to visual variations in the scene, including distractors, lighting changes, and background changes.
@@ 5 Related Work
Keypoints as Representations for Learning-based Control. Keypoints have been utilized as robust state representations for robotic manipulation in several prior works [36, 37, 38, 39]. They have been applied in areas including deformable object manipulation [40, 41], few-shot imitation learning [37], model-based reinforcement learning [39], and learning from videos [36, 42]. However, these approaches often rely on heuristic or manual keypoint selection [43, 44]. Inspired by [45], our work differs by introducing a task-driven method for automatic keypoint selection. The procedure produces flexible but expressive representations that generalize across rigid, deformable, and transparent objects, and—crucially—are robust to challenging visual disturbances.
@@ 6 Limitations and Conclusion
We present ATK, a system for automatically selecting task-relevant keypoints, learning keypoint-based policies from these representations, and successfully transferring them to the real-world evaluation scenario. Though promising, the system faces challenges in tracking and optimization. The use of 2D keypoints makes the policy sensitive to camera perspective changes, and off-the-shelf tracking modules may lack robustness for robotic applications with uncommon visual data. Additionally, the method is sensitive to hyperparameters due to the non-smooth nature of the optimization problem, making tuning difficult. Currently, the method trains different models for each individual task. A promising future direction is to incorporate a task-conditional module that can isolate task-specific keypoints and support a unified multi-task selector for diverse manipulation skills. Nevertheless, our work demonstrates the robustness of keypoint-based policies and provides an effective approach for automatic keypoint selection. Developing more automated and robust techniques to address these challenges would further extend its applicability.
@@ Acknowledgments
This work was (partially) funded by grants from the National Science Foundation NRI (#2132848), DARPA RACER (#HR0011-21-C-0171), the Office of Naval Research (#N00014-24-S-B001 and #2022-016-01 UW), and funding from the Toyota Research Institute through the University Research Program 3.0. We gratefully acknowledge gifts from Amazon, Collaborative Robotics, Cruise, and others. We would also like to thank Emma Romig, Grant Shogren for their help in setting up the hardware environments for this effort. We thank members of the Washington Embodied Intelligence and Robotics Development Lab and the Personal Robotics Lab for their thoughtful comments and feedback on versions of this draft.